3. Donald derives utility from only two goods, carrots (Qc) and donuts (Qd). (Qc and Qd are the consumption quantity of carrots and donuts). His utility function is as follows: U(Qc,Qd) = (2Qc)(Qd). Donald has an income (I) of $120 and the price of carrots (Pc) and donuts (Pd) are both $1.
a. What is Donald’s budget line?
b. What quantities of Qc and Qd will maximize Donald’s utility?
1
Expert's answer
2019-11-11T15:25:23-0500

a) Budget line


Budget line shows the combination of two products , which consumer can consume depending on his given income.

Budget line equation for Donald:

I= (Price of Carrots*Quantity of Carrots) + ((Price of Donuts*Quantity of Donuts)

"I=\lparen{P_c*Q_c}\rparen+\lparen{P_d*Q_d}\rparen"

as "P_c=P_d=1" and Income =120,

"120=\lparen{1*Q_c}\rparen+\lparen{1*Q_d}\rparen"

"\boxed{120=Qc+Qd}"


b) Utility Maximisation


Utility function of Donald: "U\lparen Q_c, Q_d \rparen = \lparen 2Q_c \rparen \lparen Q_d \rparen"

Consumer maximisation condition: "MRS= \cfrac {P_c} {P_d}"

Marginal Rate of Satisfaction (MRS) is the rate, the consumer willing to substitute one good for another. "\cfrac {P_c} {P_d}" is the price ratio of carrots and donuts.


Setting MRS equal to the price ratio:


"MRS= \frac{Marginal utility of carrots} {Marginal utility of donuts}"


"MRS= \frac{MU_c} {MU_d}" ("MU_c" and "MU_d" is remaining constant)

Then, "MU_c= 2Q_d" and "MU_d= 2Q_c"

"MRS= \frac {\not2Q_d} {\not2Q_c}"

So,

"MRS= \frac{Q_d} {Q_c}"

When setting MRS equal to price ratio: "MRS= \frac{Q_d} {Q_c} = \frac{P_d} {P_c}"

"\therefore MRS = \frac{Q_d} {Q_c} =\frac{1} {1}"

"Q_d = Q_c"


Substituting "Q_d = Q_c" in to the Budget line: "{120=Qc+Qd}"

since "Q_d = Q_c" , substitute "Q_c for Q_d"

"120=Qc+Qc"

"120=2Qc"

"\frac{120} {2}= \frac {2Qc} {2}"

"\frac{{\not120}} {\not2}= \frac {\not2Qc} {\not2}"

"\boxed{60= Q_c}"

since "Q_d = Q_c" ,

"\boxed{Q_d = 60}"

Quantities which will maximise the Donald's utility is "Q_c = 60" and "Q_d = 60" .